Given a set of data with a mean of 150 and a standard deviation of 3, what percentage of values would be greater than 159?
Answer #1

\mu =150,\sigma =3

Z=\frac{X-\mu }{\sigma }=\frac{159-150}{3}=\frac{9}{3}=3

P(X> 159)=P(Z> 3)=0.0013

  =0.13\%